OpenStudy (anonymous):

Solve the compound inequality 30 < 10u < 120. 3 < u < 4 3 > u > 12 3 > u > 4 3 < u < 12

OpenStudy (anonymous):

help

OpenStudy (souvik):

\[\frac{ 30 }{ 10 }<\frac{ 10u }{ 10 }<\frac{ 120 }{ 10 }\] or 3<u<12